OpenStudy (anonymous):

Write the equation of the line in slope-intercept form that passes through the points (-2, 6) and (-1, 4).

OpenStudy (anonymous):

use point-slope form... first get the slope of the line passing through those two points using: \(\large m=\frac{y_2-y_2}{x_2-x_1} \) then choose one of those points and plug them into point-slope form: \(\large y-y_1=m(x-x_1) \)

OpenStudy (waleed_imtiaz):

See here: slope intercept form is y= mx +c but here are two points. so first find the equation using two point form and then convert it into slope intercept form. can u do this ?

OpenStudy (anonymous):

im so confused can i just get the answer my time is running out

OpenStudy (waleed_imtiaz):

okey then I can give u the answer.... okey.... 2x+y-2=0 this is the equation when u solve for two point form. and then now convert it into slope intercept which is y= mx+c so..y=-2x+2

OpenStudy (waleed_imtiaz):

U understand it or not ?

OpenStudy (anonymous):

yea thank you very much

OpenStudy (waleed_imtiaz):

okey take care. My pleasure.